Suppose GDP is Rs.100 trillion, taxes are Rs.35 trillion, private saving is Rs.7 trillion, and public saving is Rs.4 trillion. Assuming this economy is closed, calculate



Consumption
Government Purchases
National Saving
Investment
1
Expert's answer
2021-08-31T08:44:15-0400

The consumption expenditure is calculated as follows:

C = Y - T - Private Savings

"= 100 -35 - 7"

"=58" trillion


2.

The government purchase or expenditure can be calculated as follows:

G = T - Public Savings

"=35-4"

"=31\space Trillion"

 

3.

The national savings is calculated as follows:

National Savings = Private Savings + Public Savings

"=7+4"

"=11\space Trillion"



4.

The investment is calculated as follows:

Y = C + I + G            (National Income Identity)

I = Y - C - G

"100-58-31"

"=11\space Trillion"